OpenStudy (anonymous):

How do I identify the relative maximum value for the function g(x)=2/x^2+3?

OpenStudy (anonymous):

@DemolisionWolf do you think you could help me with this? looking for the derivatives still confuses me :o

OpenStudy (anonymous):

here is a hint to work with fractions when needing to do derivatives \[\frac{ 1 }{ x^2 } = x^{-2}\]

OpenStudy (anonymous):

so does that mean that all my numbers will be negative?

OpenStudy (anonymous):

wait is it, g(x)=(2/x^2)+3 or g(x)=2/(x^2+3) ?

OpenStudy (anonymous):

the second one! im sorry, heres the problem in a picture just to make everything more clear :)

OpenStudy (anonymous):

bleh :P haha, i liked the other one more

OpenStudy (anonymous):

haha is this one harder?

OpenStudy (anonymous):

so the first step is to do the chain rule on it. ya, this one is, haha, the other one is like 2 steps.

OpenStudy (anonymous):

aw man :( lol okay, so what is the chain rule? :O

OpenStudy (anonymous):

the chain rule means we let u be equal to x^2+3 \[\frac{ d }{ du } \frac{ 1 }{ u }\frac{ du }{ dx }\] does this sound familiar?

OpenStudy (anonymous):

not really :o

OpenStudy (anonymous):

does thhat mean that i add x^2+3 to the top equation?

OpenStudy (anonymous):

hmmm... I wonder if there is another way to solve this w/o the chain rule...

OpenStudy (anonymous):

hm, well I could definitely try it with the chain rule if you thinks its the best way

OpenStudy (anonymous):

i'm not sure how else to solve it.. and getting into the chain rule is a bit longer than I had wanted to spend on this, so we will go quick, cuz I gotta go at 5:30 ^_^

OpenStudy (anonymous):

oh okay, i totally understand :)

OpenStudy (anonymous):